OpenStudy (anonymous):

PLEASE HELP ME!! Medal and Fanning! For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10

OpenStudy (anonymous):

2y + 10 2y + 10 2y - 10 -2y - 10

OpenStudy (the_fizicx99):

After you have solved for "x" you would plug in what you have solved for into the other equation in the x's. I don't know if that makes sense 2x + y = 4 3x + 2y = 4 2x + y = 4 -y -y 2x = -y + 4 /2 /2 /2 x = -(y/2) + 2 now plug that into 3x + 2y = 4 3(-(y/2) + 2) + 2y = 4 and then you would just solve.

OpenStudy (anonymous):

So would it be -2y+10?

OpenStudy (the_fizicx99):

IF you isolated x in the equation: -x - 2y = -10 You would get; -x - 2y = -10 +2y +2y -x = 2y - 10 /-1 /-1 /-1 x = -2y + 10 Now plug in that \(\ -2y + 10 \) into 3x + y = 8 and to do that, simply plug in \(\ \color{green}{-2y + 10}\) into \(\ 3(\color{green}{???}) + y = 8 \)

OpenStudy (anonymous):

Thanks soo much @tHe_FiZiCx99 !!!! =^.^=

OpenStudy (the_fizicx99):

You're welcome :P

OpenStudy (anonymous):

:)

OpenStudy (hahawowoh):

so what was the answer featheres